OpenStudy (anonymous):

find the points on the curve y=4x^3 -6x^2- 24x +30 where the tangent is parallel to the x axis. please help and show all steps!

OpenStudy (turingtest):

"tangent is parallel to the x axis" means when the derivative of f(x)=0 i.e. solve y=f(x)=4x^3 -6x^2- 24x +30 dy/dx=f'(x)=12x^2-12x-24=0

OpenStudy (anonymous):

so are the points (-1, 44) and (2, -10)?